### v1

# The Preheating Stage on The Starobinsky Inflation after ACT

Abstract
In this paper, we reinvestigate the Starobinsky inflation and its reheating features after the ACT results. To make the Starobinsky model satisfy ACT with $68\%$ CL, the number of e-folds increases to $>70$ while the reheating temperature drops to $10^{4}$ GeV. The preheating stages and the reheating temperature need to be adjusted drastically in favor of the lower temperature. We obtain that the preheating stage of Starobinsky inflation should be mixed between tachyonic and non-tachyonic preheating to favor the analytical estimation, PBH abundance, and the low reheating temperature at once. For that purpose, a varied momentum is necessary to obtain the favored results.
1 Introduction
The Starobinsky model [1], as the simplest example of a modified gravity theory [2, 3, 4], is considered one of the most successful inflationary models. Its simplicity lies in the fact that it relies solely on gravity as the driving mechanism for inflation. Despite this simplicity, it remains consistent with the Cosmic Microwave Background (CMB) observations from Planck data [5]. In recent years, this model and its various extensions have continued to attract considerable attention and discussion [6, 7, 8, 9, 10, 11, 12]. However, the Starobinsky model is now challenged by newly established observational data from the Atacama Cosmology Telescope (ACT) [13]. This new data shows that the slow-roll parameter is $n_{s}=0.9709\pm 0.0038$. When combined with measurements from CMB lensing, Baryon Acoustic Oscillation (BAO), and the Dark Energy Spectroscopic Instrument (DESI), the extended result yields $n_{s}=0.9743\pm 0.0034$. This result indicates that the Starobinsky model is excluded at the $2\sigma$ level for $N\approx 60$. The model remains viable if a large number of e-folds ($N\approx 75$) is allowed, which leads to a lower reheating temperature [14, 15, 16]. Alternatively, this issue can be addressed by introducing higher-curvature perturbations to satisfy the new constraint [17].
The reheating temperature of the Starobinsky model is found to be relatively low $\sim 10^{9}$ GeV [2, 18] compared to other models, such as Higgs inflation [19]. This lower reheating temperature is often preferred in inflationary cosmology, as it helps avoid the overproduction of unwanted relics [20]. For example, Refs. [21, 22] argue [^1] See also Ref. [23]. that the reheating temperature should be below $10^{9}$ GeV to prevent excessive production of gravitino. Such a low reheating temperature significantly influences the thermal history of the universe, particularly with respect to Baryogenesis [24] and the production of dark matter. However, as for leptogenesis, it requires at least several MeV of reheating temperatures to operate successfully [25, 26]. In summary, estimation of the reheating temperature range is wide, from a few MeV up to nearly the Planck scale. This broad range suggests that the reheating temperature is still poorly constrained. In addition to the ACT observation results, the Starobinsky inflation model is expected to yield a much lower reheating temperature.
The preheating stage of the Starobinsky model has received comparatively less attention in the literature than its inflationary phase. This is largely because the inflationary dynamics are better constrained, whereas the preheating phase tends to be more model-dependent [27, 28]. In the Starobinsky model, preheating is primarily driven by the interaction of the produced particles with gravity. While this makes it a simple and potentially promising scenario, our limited understanding of gravity makes it more challenging to analyze fully. In this paper, we investigate the preheating mechanism in the Starobinsky model, focusing on the efficient energy drain by preheating, the constraints from primordial black hole (PBH) production, the duration of the preheating stage, and the resulting reheating temperature.
This paper is organized as follows: In Section 2, we discuss the inflationary description of the Starobinsky model. In Section 3, we examine the preheating phase of the Starobinsky model, which is non-minimally coupled to scalars. In Section 4, we estimate the duration of preheating and the resulting reheating constraints. In Section 5, we consider the possibility of fermionic preheating. We also discuss primordial black hole (PBH) formation in Section 6 and the ceased-out mechanism in Section 7. The reheating scenario is discussed in Section 8. Finally, we present our conclusions in Section 9.
2 The inflationary model
We start this section by writing the action in the Jordan frame as [^2] In this paper, we used (,+,+,+)(-,+,+,+)( - , + , + , + ) convention.
[…]
where $g_{J}$ is the determinant of the metric tensor $g^{\mu\nu}_{J}$, $M_{p}=1/\sqrt{8\pi G}$ is the reduced Planck mass, $R_{J}$ is the Ricci scalar, and $M$ is the effective inflaton’s mass at the end of inflation. The subscript $J$ represents the Jordan frame. By using the conformal transformation as
$g^{\mu\nu}_{E}\equiv\Omega^{-2}g^{\mu\nu}_{J},\hskip 28.45274pt\Omega^{2}\equiv\frac{df(R_{J})}{dR_{J}}=1+\frac{R_{J}}{3M^{2}}\equiv\exp\left({\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)$ (2.2)
We obtain the action in the Einstein frame as follows
$S_{E}=\int d^{4}x\sqrt{-g_{E}}\left[\frac{1}{2}M_{p}^{2}R_{E}-\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)\right],$ (2.3)
where ${R}_{E}=R_{J}\Omega^{-2}-6\Omega^{-3}\square\Omega$ and
$V(\phi)=\frac{3}{4}M_{p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)^{2}.$ (2.4)
For the rest of this paper, we will omit the subscript $E$ for simplicity.
Firstly, we will write the scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, slow-roll parameters ($\epsilon$ and $\eta$) and the scalar power spectrum ($A_{s}$) as
$n_{s}=1-6\epsilon+2\eta,\quad r=16\epsilon,\quad A_{s}=\frac{V(\phi)}{24\pi^{2}\epsilon}\quad\epsilon=\frac{1}{2}M_{p}^{2}\left(\frac{V^{\prime}}{V}\right)^{2}\quad\eta=M_{p}^{2}\frac{V^{\prime\prime}}{V}.$ (2.5)
The prime denotes the derivative with respect to scalaron $\phi$. By using the ACT observation result [13] as $n_{s}=0.9743\pm 0.0038$, $A_{s}\approx 2.13\times 10^{-9}$ at the pivot scale $\kappa_{*}=0.05$ Mpc-1, we obtain the new constraint on the $\phi_{\text{ini}}$, corresponds to the field value of the scalaron at $\kappa_{*}$, to be
$\phi_{\text{ini}}\approx 5.72\hskip 2.84526ptM_{p}.$ (2.6)
The inflaton’s field value at the end of inflation is obtained by setting $\epsilon=1$, which corresponds to the violation of the slow-roll to be
$\phi_{\text{end}}\simeq 0.94\hskip 2.84526ptM_{p}.$ (2.7)
By using the last two results, we can obtain the number of e-folds $N$ as
$N=\frac{1}{M_{p}^{2}}\int^{\phi_{\text{ini}}}_{\phi_{\text{end}}}\frac{V}{V^{\prime}}\simeq 75.51,$ (2.8)
much higher than the previously predicted. [^3] One should also note that slight changes in the constraint of $n_{s}$ could potentially alter the results of other parameters greatly. This large number of e-folds is expected for a large field value of the inflaton. [^4] The chosen e-fold greater than 60 is rather implausible but still acceptable [29]. In addition, the effective mass of the scalaron $M$ is also affected by the new ACT constraint on the scalar power spectrum $A_{s}$ as
$M\simeq 9.07\times 10^{-6}M_{p},$ (2.9)
which is slightly lower than the Planck result [5].
3 The preheating
When $\phi\lesssim M_{P}$, the inflaton oscillates, marking the end of inflation. The particle produced by this mode is non-perturbative. If we assume $\chi$ is dominantly produced during this mode, which is presented by the action:
$S_{\chi}=\int d^{4}x\sqrt{-g_{J}}\left[\frac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\frac{1}{2}m_{\chi}^{2}\chi^{2}-\xi R_{J}\chi^{2}\right],$ (3.1)
we can evaluate its particle production. In this paper, we set $\xi=1$ in both analytical and numerical calculations to avoid the naturalness issue. The equation of motions (e.o.m) in the Einstein frame from varying $\delta\phi$ and $\delta\chi$ respectively are
$\begin{split}&\ddot{\phi}_{k}+3H\dot{\phi}_{k}+\sqrt{\frac{3}{2}}\frac{1}{M_{p}}e^{-3\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\left(\dot{\chi}^{2}_{k}+\frac{k_{\chi}^{2}}{a^{2}}\chi^{2}_{k}\right)-\sqrt{\frac{2}{3}}e^{-2\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\frac{m_{\chi}^{2}\chi^{2}_{k}}{M_{p}}+\frac{k^{2}_{\phi}}{a^{2}}\phi_{k}\\ &\hskip 5.69046pt+\sqrt{\frac{3}{2}}M_{p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\right)\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)+\sqrt{\frac{3}{2}}\frac{\xi M^{2}}{M_{p}}e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\left(2e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}-1\right)\chi^{2}_{k}=0\end{split}$ (3.2)
and
$\begin{split}&\ddot{\chi}_{k}+3H\dot{\chi}_{k}-3\sqrt{\frac{2}{3}}\frac{\dot{\phi_{k}}\dot{\chi_{k}}}{M_{p}}\\ &\hskip 28.45274pt+\left[\frac{k^{2}_{\chi}}{a^{2}}-3\sqrt{\frac{2}{3}}\frac{k_{\phi}k_{\chi}}{a^{2}}\frac{\phi_{k}}{M_{p}}+e^{\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}m_{\chi}^{2}-3\xi M^{2}e^{2\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}-1\right)\right]\chi_{k}=0,\end{split}$ (3.3)
where we used the Heisenberg representation as
$\begin{split}&\phi(\textbf{x},t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}\left(\hat{a}_{k}\phi_{k}(t)e^{-i\textbf{k}_{\phi}\cdot\textbf{x}}+\hat{a}^{\dagger}_{k}\phi^{*}_{k}(t)e^{i\textbf{k}_{\phi}\cdot\textbf{x}}\right),\\ &\chi(\textbf{x},t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}\left(\hat{a}_{k}\chi_{k}(t)e^{-i\textbf{k}_{\chi}\cdot\textbf{x}}+\hat{a}^{\dagger}_{k}\chi^{*}_{k}(t)e^{i\textbf{k}_{\chi}\cdot\textbf{x}}\right),\end{split}$ (3.4)
on calculating Eq. (3.2) and (3.3). At the end of inflation, Eq. (3.3) can be approximated by using $\phi=\tilde{\phi}\sin(Mt)$ and $a=1$ when $\tilde{\phi}<M_{p}$ as
$\ddot{\chi}_{k}+\Bigg{[}\left(k^{2}_{\chi}+m_{\chi}^{2}\right)+\left(\sqrt{\frac{2}{3}}\frac{m_{\chi}^{2}+3\xi M^{2}}{M_{p}}-3\sqrt{\frac{2}{3}}\frac{k_{\phi}k_{\chi}}{M_{p}}\right)\tilde{\phi}\sin(Mt)\Bigg{]}\chi_{k}\simeq 0.$ (3.5)
For the near-zero crossing ($\sin(Mt)\simeq Mt$), we can approximate Eq. (3.5) as
$\frac{d^{2}{\chi}_{k}}{dq^{2}}+\left(p^{2}+q\right)\chi_{k}=0,$ (3.6)
where
$\begin{split}&p^{2}=\left(k_{\chi}^{2}+m_{\chi}^{2}\right)\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{-2/3}\\ &q=\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{1/3}t.\end{split}$ (3.7)
This simplification will help us to understand the numerical results depicted in Fig. 1, which are obtained by solving Eqs. (3.2) and (3.3). In the following, we will describe two cases in which the fine-tuning of $k_{\phi}$ and $k_{\chi}$ in Fig. 1 could affect the preheating stage.
Figure 1: The preheating of $\phi_{k}$ oscillation (black-dashed line) and $\chi_{k}$ (red-solid line) by momentum-dependent $k_{\phi}$ and $k_{\chi}$. The figure is based on the numerical calculation of (3.2) and (3.3). For this numerical result, we used $\xi=1$ and $M=9.07\times 10^{-6}M_{p}$, while $M_{p}=1$. In the left figure, we used $m_{\chi}=10^{-6}M_{p}$, $k_{\phi}=10^{-7}M_{p}$ and $k_{\chi}=10^{-4}M_{p}$. In the right figure, we used $m_{\chi}=10^{-6}M_{p}$, $k_{\phi}=10^{-3.5}M_{p}$ and $k_{\chi}=10^{-4}M_{p}$. In our numerical calculation, we set $\phi_{k}(0)=M_{p}$, $\frac{d\phi_{k}}{dt}\big{|}_{t=0}=0$, $\chi_{k}(0)=0$, and $\frac{d\chi_{k}}{dt}\big{|}_{t=0}=0$. The vertical axis corresponds to $\phi_{k}$ and $\chi_{k}$ in the unit of $M_{p}$. The horizontal axis corresponds to $t$ in the unit of $\pi M^{-1}$.
  • Case 1
    As a starting point, the left panel of Fig. 1 shows that the energy transfer from the inflaton field $\phi$ to the particle production of $\chi$ is inefficient. This condition is satisfied analytically if (see Eqs. (3.5) and (3.6))
    $m_{\chi}^{2}+3\xi M^{2}\gg 3k_{\phi}k_{\chi}.$ (3.8)
    In the conventional theory of inflation, the inflaton should have vanished in the early universe as it decayed into smaller particles. This ensures that there is no inflaton today. However, Starobinsky inflation is rather special. As this inflation is gravity-driven, after the reheating process, it allows the scalaron (or gravity) to remain if all decay channels do not exist. Consequently, $m_{\chi}$ does not need to be smaller than $M$. As we will shortly see, however, these conditions are disfavored.
    Suppose the condition $2m_{\chi}>M$ holds, the perturbative decay of the scalaron to $\chi$ is kinematically blocked. In this case, the production of $\chi$ can only occur via parametric resonance during preheating. However, since this form of preheating is extremely inefficient (see the left panel of Fig. 1), it implies that, by the end of the preheating stage, the scalaron would predominantly remain without any viable decay channels. As a consequence, the universe would lack a sufficient number of particles, contradicting current observational data. Therefore, the condition $2m_{\chi}>M$ is not a realistic scenario.
    On the other hand, if $2m_{\chi}<M$, the perturbative decay is kinematically allowed. Thus, the reheating temperature can be evaluated from this decay, although it can also be determined from the decay of the daughter fields of $\chi$, which is equally possible. We will discuss this case in the section 8.
    For $p^{2}\ll q$, we can estimate the particle production of $\chi$ by the single crossing of $\phi$ as [19, 30]
    $\delta\rho=\int^{\infty}_{0}\frac{d^{3}k_{\chi}}{(2\pi)^{3}}e^{-\pi p^{2}}\bar{m}$ (3.9)
    where
    $\begin{split}&\bar{m}\simeq\sqrt{\sqrt{6}|k_{\phi}k_{\chi}|\frac{\tilde{\phi}}{M_{P}}}\end{split}.$ (3.10)
    Analytically, by calculating Eq. (3.9), we obtain that in the first zero crossing (after the end of inflation), it produces
    $\delta\rho_{\text{(Case 1)}}\simeq 0.26\times\xi^{3/2}M^{4}\sim 10^{-21}\xi^{3/2}M_{p}^{4}.$ (3.11)
    In this calculation, to achieve the same energy as inflaton ($\frac{1}{2}M^{2}\tilde{\phi}^{2}\sim 10^{-11}M_{p}^{4}$), we require $\xi\gg\mathcal{O}(1)$, which is strongly disfavored due to naturalness issues. We conclude that preheating in Case 1 is generally inefficient. We will see later in Section 4 that applying only this preheating mode alone in Starobinsky inflation is strongly disfavored.
  • Case 2
    The tachyonic preheating could also appear due to the momentum of both $\phi$ and $\chi$ as
    $m^{2}_{\chi}+3\xi M^{2}\ll 3k_{\phi}k_{\chi},$ (3.12)
    which later we called by the tachyonic conditions. The growth of particle production due to tachyonic preheating evolves as $\propto e^{2\mu_{p}Mt}$. Thus, it is easier to observe the effect of tachyonic preheating in the right panel of Fig. 1. One can see in the figure that, as we set the value of $k_{\phi}k_{\chi}$ slightly higher than $m^{2}_{\chi}+3\xi M^{2}$, the growth of $\chi$ becomes extremely large, and the amplitude of $\phi$ drops significantly after the first crossing.
Based on the two cases above, it is clear that we can understand the two possibilities of the preheating process. Additionally, Eq. (3.5) can be redefined by
$\begin{split}&Mt\equiv 2z-\pi/2,\hskip 28.45274ptA\equiv 4\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{M^{2}}\right),\\ &Q\equiv\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}.\end{split}$ (3.13)
So we obtain the famous Mathieu equation as
$\frac{d^{2}\chi_{k}}{dz^{2}}+(A-2Q\cos(2z))\chi_{k}=0.$ (3.14)
The numerical calculation of the Mathieu instability is shown in Fig. 2. The growth of the $\chi_{k}$ is depicted by the characteristic exponent $\mu_{p}$, which can be written analytically as
$\mu_{p}=\frac{1}{2\pi}\ln\left(1+2e^{-\pi p^{2}}-2\sin\theta_{\mu}e^{-\frac{\pi p^{2}}{2}}\sqrt{1+2e^{-\pi p^{2}}}\right),$ (3.15)
where $\theta_{\mu}$ is a phase. It is shown that in Case 1, where the preheating is narrow, it should lie in the stability region, while Case 2 should belong to the instability region (see Fig. 2).
Figure 2: Mathieu instability region related to the semi-analytical result depicted in Eq. (3.14). The right legend corresponds to the value of $\mu_{p}$. Note that $Q$ can have negative values in our model. The white region corresponds to the stable condition.
4 The preheating duration and reheating
Generally, the preheating duration plays a role in determining the reheating temperature. If the duration is short, the reheating temperature is high, and vice versa [30]. In that case, a proper preheating duration is necessary for a viable reheating temperature. However, it is still model-dependent. Some peculiar models with long durations can also yield high reheating temperatures [^5] Higgs inflation could be one of the best examples. Due to the large non-minimal coupling, the duration of preheating should be long [30]. But its instant-type decay from gauge bosons to fermions gives a high reheating temperature, $T_{\text{reh}}\sim 10^{14}$ GeV [19].
To obtain the preheating duration, we will consider Eq. (2.3) and write the (oversimplified) e.o.m. together with the first and second Friedmann equations as [^6] The e.o.m. and two Friedmann equations here are rather oversimplified. As the purpose is to estimate the order of the time $t$, we think it is sufficient.
$\begin{split}&\ddot{\phi}+3H\dot{\phi}+M^{2}\phi=0,\\ &H^{2}=\frac{1}{3M_{p}^{2}}\left(\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}M^{2}\phi^{2}\right),\\ &\dot{H}=-\frac{\dot{\phi}^{2}}{2M_{p}^{2}}.\end{split}$ (4.1)
If we define $\dot{\phi}=\sqrt{6}HM_{p}\cos\theta$ and $M\phi=\sqrt{6}HM_{p}\sin\theta$, straightforwardly, we obtain
$\tilde{\phi}=2\sqrt{\frac{2}{3}}\frac{M_{p}}{Mt}.$ (4.2)
Lastly, following the method as Ref.[30], when the Jordan frame coincides with the Einstein frame, we may set $\tilde{\phi}_{\text{end}}=M_{p}$ and $\tilde{\phi}_{\text{reh}}=M$, yielding [31, 30]
$\Delta t=|t_{\text{reh}}-t_{\text{end}}|\simeq 2\sqrt{\frac{2}{3}}\frac{M_{p}}{M^{2}}\sim 10^{5}\pi M^{-1},$ (4.3)
where $t_{\text{reh}}$ is the time when the radiation-dominated era starts and $t_{\text{end}}$ is the time when preheating starts after the end of inflation. If we consider the result in the left figure of Fig. 1, the broad resonance still occurs at $t\gg 10^{5}\pi M^{-1}$. In that case, we conclude that such a mechanism is disfavored if preheating only adopts the Case 1-type. On the other hand, if the preheating only adopts Case 2, the preheating duration is too short compared to our analytical result in Eq. (4.3), so this is also disfavored. But realistically, during preheating, both $k_{\phi}$ and $k_{\chi}$ may not be fixed [30]. In this context, the dominant preheating process may arise from Case 1 being interrupted by Case 2 during the early or late stages of preheating. This combination allows for a preheating duration that aligns with the analytical result presented in Eq. (4.3). As will be discussed in Section 7, we will identify the specific time of the preheating period during which tachyonic preheating intervenes. [^7] In fact, we will get an indication that tachyonic preheating occurs during the initial stage of preheating from the discussion on PBH formation (see Section 6). To estimate the reheating temperature, it is necessary to write the formula as
$\frac{\kappa_{*}}{a_{0}H_{0}}=\frac{a_{\kappa}}{a_{\text{end}}}\frac{a_{\text{end}}}{a_{\text{reh}}}\frac{a_{\text{reh}}}{a_{\text{eq}}}\frac{a_{\text{eq}}}{a_{0}}\frac{H_{\text{eq}}}{H_{0}}\frac{H_{\kappa}}{H_{\text{eq}}},$ (4.4)
where $a_{0}=1$ and $H_{0}=66.78\hskip 2.84526pt\text{km Mpc}^{-1}\text{s}^{-1}$ [13] represent the scale factor and Hubble parameter today. $a_{\kappa}$ and $H_{\kappa}$ represent the scale factor and Hubble parameter at the pivot scale $\kappa_{*}=0.05\hskip 2.84526pt\text{Mpc}^{-1}$. The subscripts ’eq’, ’end’, and ’reh’ correspond to ’matter-radiation equality’, ’end of inflation’, and ’reheating’, respectively. Note that we solve Eq. (4.4) by using $T_{\text{eq}}=0.8$ eV, and the end of preheating is evaluated at $\tilde{\phi}=M$. Finally, by using (4.4) with e-fold $N=75.5$, we obtain an upper bound for the reheating temperature as $T_{\text{reh}}\sim 10^{4}$ GeV. This value is five orders of magnitude lower than the previous prediction [2].
5 On the fermionic preheating
In this part, we will discuss the possibility that, instead of the scalar $\chi$, the scalaron is coupled with fermions. So far, we have only discussed the bosonic preheating, even though fermionic preheating is equally possible. This type of preheating has already been discussed in, e.g., Refs. [32, 33, 34, 35]. In this section, we will qualitatively discuss the possibility of this type of preheating.
If the scalaron is only coupled with certain fermions, at a large coupling [33], the preheating is similar to the bosonic case. However, it is heavily suppressed by the Pauli exclusion principle, in contrast to the Bose-Einstein condensate (BEC) that can occur during bosonic preheating. In this case, the inflaton energy transfer during resonance is mostly converted into the expansion of the universe. Fermion production becomes less suppressed once the universe expands. However, even after the end of preheating, the potential of the scalaron remains quadratic ($\propto\phi^{2}$), which still favors a matter-dominated regime. Fermions are more likely to be produced at the end of the preheating phase due to the looseness of the Pauli blocking.
Even if tachyonic preheating occurs, fermionic particle production remains heavily suppressed, unlike in the bosonic case. This means fermionic preheating is less efficient, even compared to Case 1 of bosonic preheating. Consequently, the duration of preheating is expected to exceed our analytical prediction. Finally, we safely said that preheating via this mode is strongly disfavored.
6 The PBH formation from preheating
In this section, we will discuss PBH formation due to preheating in the Starobinsky model. We will show later that the PBH formation is consistent with it being produced during the early stage of preheating.
Straightforwardly, the constraint on the PBH can be written as
[…]
where $P(\delta)$ represents the probability distribution function, and it can be written as
$P(\delta)=\frac{1}{\sqrt{2\pi}\sigma_{\delta}}\exp\left(-\frac{\delta^{2}}{2\sigma_{\delta}^{2}}\right).$ (6.2)
By using $\beta_{\text{PBH}}=10^{-20}$ and $\delta=\delta_{c}=0.7$ [36, 37] on Eq. (6.1), we obtain the mass variance $\sigma_{\delta}$ as
$\sigma_{\delta}\simeq 0.08.$ (6.3)
However, we need to skip this part before putting the constraint into our parameters.
For the production of PBH during preheating, the strong parametric resonance should be applied. Our assumptions implied that PBH is most likely produced when [38, 36]
$|Q|\equiv\left|\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}\right|\gg p^{2}.$ (6.4)
The power spectrum of the variable $x$ can be expressed as [39]
$\mathcal{P}_{(x)}=\frac{k_{x}^{3}}{2\pi^{2}}\braket{}{x_{k}}{{}^{3}},$ (6.5)
where $k_{x}=|\textbf{k}_{x}|$ represents the comoving wavenumber associated with $x$. Note that $x_{k}$ denotes the inverse Fourier transform of $x$. The influence of preheating on $\chi$ production can be determined as [40, 38, 36]
$\mathcal{P}_{\delta\chi}=\mathcal{P}_{\delta\chi}|_{\text{end}}\exp\left(2\mu_{p}M(t-t_{\text{end}})\right).$ (6.6)
The characteristic exponent $\mu_{p}$ here is evaluated at $\theta_{\mu}=0$, in which $\mu_{p}=(2\pi)^{-1}\ln(1+2\exp(-\pi p^{2}))$, where $p$ is taken from Eq. (3.7) as
$p^{2}=\frac{k_{\chi}^{2}+m_{\chi}^{2}}{\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{2/3}}=\frac{1}{18\sqrt{Q}}\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{k_{\text{end}}^{2}}\right),$ (6.7)
where we introduce
$k_{\text{end}}^{2}=\frac{1}{18\sqrt{2}}\left|\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}\right]^{1/6}\right|M^{5/3}$ (6.8)
as the comoving wavenumber evaluated at the Hubble exit at the end of inflation. For $Q\gg 1$, in which it is required to get the largest production, we get $\mu_{p}=\mu_{0}\simeq\frac{1}{2\pi}\ln 3$ and our chosen parameters should be tachyonic ($m_{\chi}^{2}+3\xi M^{2}\ll 3k_{\phi}k_{\chi}$). In addition, we will calculate the non-adiabatic power spectrum of $\chi$ as [38, 36]
$\mathcal{P}_{\zeta_{\text{n-ad}}}(k_{\chi})\simeq\mathcal{A}\left(\frac{k_{\chi}}{k_{\text{end}}}\right)^{3}I(p,M\Delta t),$ (6.9)
where [^8] The third bracket is coming from the interaction coupling between $\phi$ and $\chi$ via $\mathcal{L}_{\text{int}}=\sqrt{\frac{2}{3}}\frac{m^{2}_{\chi}}{M_{P}}\phi\chi^{2}$ divided by mass scale $M$.
$\mathcal{A}\equiv\frac{3\times 2^{3/2}}{\pi^{6}\mu_{p}^{2}}\left(\frac{\tilde{\phi}}{M_{p}}\right)^{2}\left(\frac{H_{\text{end}}}{M}\right)^{4}\left(\sqrt{\frac{2}{3}}\frac{m^{2}_{\chi}}{MM_{P}}\right)^{2}Q^{-1/4}$ (6.10)
and
[…]
Note that $\vartheta$ is the angle between $p$ and $p^{\prime}$. To solve the last equation, we need to impose [36]
$\mu_{p^{\prime}}\simeq\mu_{p-p^{\prime}}\approx\mu_{0}+\frac{p^{\prime 2}}{3}.$ (6.12)
So we obtain
[…]
Finally, the mass variance $\sigma_{\delta}$ during the matter-dominated era can be calculated as follows [41]
$\sigma_{\delta}^{2}=\frac{16}{81}\int^{\infty}_{0}W^{2}(\tilde{k}_{\chi},\mathcal{R})\left(\frac{\tilde{k}_{\chi}}{{k}_{\chi}}\right)^{4}\mathcal{P}_{\zeta_{\text{n-ad}}}\frac{d\tilde{k}_{\chi}}{\tilde{k}_{\chi}},$ (6.14)
where we introduced the window function $W(\tilde{k}_{\chi},\mathcal{R})=\exp\left(-\tilde{k}_{\chi}^{2}\mathcal{R}^{2}/2\right)$, with $\mathcal{R}=1/k_{\psi}$. Finally, we can use $\sigma_{\delta}$ calculated from Eqs.(6.14), (6.1) and (6.3) to obtain the constraint on PBH production to be
$5.43\times 10^{-3}\gtrsim\frac{e^{4\mu_{0}\cdot M\Delta t}}{(M\Delta t)^{3/2}}\cdot\frac{m^{4}_{\chi}}{M^{4}M_{p}^{2}}\frac{k^{\frac{5}{2}}_{\chi}}{\sqrt{k_{\phi}}}.$ (6.15)
Eq. (6.15) shows the constraints on $k_{\chi}$, $k_{\phi}$, $m_{\chi}$, and $\Delta t$ simultaneously. In the following, we describe the numerical results for Eq. (6.15) in Table 1. In this table, we set the parameters based on Case 2. The preheating duration is analytically found to be $\sim 10^{5}\pi M^{-1}$, and it is shown that a sufficient amount of PBHs is already produced at the early stage of preheating. However, due to the rapidly expanding universe, the overdense region is wiped out very quickly, leading to the rapid suppression of PBH formation as well [36].
$m_{\chi}$ $k_{\phi}$ $k_{\chi}$ $\Delta t$ ($\pi M^{-1}$)
$10^{-12}M_{p}$ $10^{-3.5}M_{p}$ $10^{-4}M_{p}$ $382$
$10^{-12}M_{p}$ $10^{-4}M_{p}$ $10^{-4}M_{p}$ $380$
$10^{-12}M_{p}$ $10^{-5}M_{p}$ $10^{-4}M_{p}$ $377$
$10^{-12}M_{p}$ $10^{-6}M_{p}$ $10^{-4}M_{p}$ $374$
$10^{-12}M_{p}$ $10^{-7}M_{p}$ $10^{-4}M_{p}$ $364$
$10^{-12}M_{p}$ $10^{-8}M_{p}$ $10^{-4}M_{p}$ $359$
Table 1: The duration of the required PBH abundance. The chosen parameters are assumed to follow the tachyonic conditions.
In our calculation of Eq. (6.15), the duration range to obtain a sufficient amount of PBH is always $\sim 10^{2}\pi M^{-1}$, insensitive to the chosen parameters of $m_{\chi}$, $k_{\phi}$, and $k_{\chi}$. In that case, the PBH constraint may not jeopardize our model.
7 The ceased-out mechanism
In this section, we will discuss the cease-out mechanism based on Ref. [30]. In this mechanism, the momenta $k_{\phi}$ and $k_{\chi}$ are not fixed; instead, they become time-dependent. As both momenta become functions of time, an appropriate functional form can be imposed to obtain the desired model.
In our previous calculation, it was shown that preheating in this model can be estimated to end at $\sim 10^{5}\pi M^{-1}$. Both cases indicate that neither mechanism alone can satisfy the prediction. Thus, the most plausible approach is to combine both mechanisms into one. In doing so, either $k_{\phi}$ or $k_{\chi}$ (or both) must be time-dependent.
In the numerical calculation of the equations (3.2) and (3.3), both $k_{\phi}$ and $k_{\chi}$ are fixed. However, in the cease-out mechanism, we assume $k_{\phi}$ to be a function of $t$ as
$k_{\phi}\equiv k^{0}_{\phi}e^{\alpha t},$ (7.1)
where $k^{0}_{\phi}$ is the initial amplitude of $k_{\phi}$ and $\alpha$ is a constant. As the production of $\chi$, especially during the first stage of preheating, should be heavy non-relativistic, it is implied that the momentum $k_{\chi}$ should be small.
In the left panel of Fig. 3, we show the condition when the ceased-out mechanism is imposed on $k_{\phi}$. In this numerical calculation, we set $\alpha=-10^{-5}$ in units of $M\pi^{-1}$. In this figure, $k_{\phi}$ is initially at $k_{\phi}^{0}=10^{-3.5}M_{p}$ but decreases exponentially. Consequently, the tachyonic growth stops after $t$ reaches $\sim 50000$ (in $\pi M^{-1}$). Thus, the time when preheating stops is favorable for our analytical prediction. On the other hand, in the right figure of Fig. 3, if we set $\alpha=10^{-5}M\pi^{-1}$ with initial momentum $k^{0}_{\phi}=10^{-5}M_{p}$, preheating is initially inefficient but becomes more efficient due to tachyonic preheating. At $t\sim 10^{5}$ (in $\pi M^{-1}$), preheating becomes narrow. However, to match the previous discussion about PBH, the left figure of Fig. 3 is disfavored. PBH formation should be significant in the early stage of preheating. Thus, a large momentum $k_{\phi}$ is necessary at the beginning of preheating.
Figure 3: The preheating of $\phi_{k}$ (black-dashed line) and $\chi_{k}$ (red-solid line) oscillations by momentum-dependent $k_{\phi}$ and $k_{\chi}$ affected by the ceased-out mechanism. In this numerical calculation of the left panel, we imposed $m_{\chi}=10^{-6}M_{p}$, $M=9.7\times 10^{-6}M_{p}$, $M_{p}=1$, $\xi=1$, $k_{\chi}=10^{-4}M_{p}$ and $k_{\phi}$ is written as Eq. (7.1) with $\alpha=-10^{-5}M\pi^{-1}$ and $k^{0}_{\phi}=10^{-3.5}M_{p}$. In the right panel, we used $k_{\phi}^{0}=10^{-5}M_{p}$ and $\alpha=10^{-5}M\pi^{-1}$; otherwise, all remained parameters are the same. In our numerical calculation, we also set $\phi_{k}(0)=M_{p}$, $\frac{d\phi_{k}}{dt}\big{|}_{t=0}=0$, $\chi_{k}(0)=0$, and $\frac{d\chi_{k}}{dt}\big{|}_{t=0}=0$.
8 Reheating scenario
In this section, we will discuss the possible reheating scenario in this model. It is noted that the reheating stage is the condition where the cosmological model can be closely linked to SM particle physics. However, it turns out that calculating the reheating temperature is not straightforward. We will see in the following subsection that several mechanisms can appear, and we will discuss which mechanism is the most favored.
8.1 The Perturbative decay of the scalaron to scalar $\chi$
The most traditional depiction of the reheating temperature is based on the simple perturbative decay of the inflaton into light and relativistic particles. It is assumed that the energy of these light particles is converted into radiation. If we consider the scalaron decaying to $\chi$ via the Lagrangian
$\sqrt{\frac{3}{2}}\frac{\phi}{M_{p}}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi,$ (8.1)
the decay rate, within approximation $m_{\chi}\ll M$, can be written as
$\Gamma(\phi\rightarrow\chi\chi)\simeq\frac{3}{128\pi}\frac{M^{3}}{M_{p}^{2}},$ (8.2)
and the reheating temperature can be easily evaluated as
$T_{\text{reh}}=\left(\frac{90}{g^{*}\pi^{2}}\right)^{1/4}\sqrt{M_{p}\Gamma},$ (8.3)
where $g^{*}\sim 100$ is the SM degrees of freedom. As $M$ is well-constrained by ACT, it is clear that the reheating temperature can be obtained as
$T_{\text{reh}}\approx 3.14\times 10^{9}\text{GeV}.$ (8.4)
However, this value is 5 orders of magnitude larger than our prediction after applying the ACT constraint. In this case, the reheating temperature from this perturbative mode is strongly discouraged. One should note that this value could be enhanced by BEC effects, which may increase the reheating temperature.
8.2 The Perturbative decay of the scalaron to fermion $\psi$
As previously remarked, the perturbative decay of the scalaron to $\chi$ is heavily suppressed by ACT results. In that case, the interaction between the scalaron and fermions is described by the Dirac Lagrangian in the Einstein frame, which reads
$\mathcal{L}_{\text{Dirac}}=e^{-3\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m_{\psi})\psi,$ (8.5)
where $\bar{\psi}$ and $\psi$ are antifermion–fermion pairs, $\gamma^{\mu}$ is the Dirac gamma matrix, and $m_{\psi}$ is the mass of the fermions. Due to the kinetic term of the Dirac Lagrangian being on-shell, the interaction term of the scalaron with $\bar{\psi}\psi$ is controlled by
$3\sqrt{\frac{2}{3}}\frac{m_{\psi}}{M_{p}}\phi\bar{\psi}\psi.$ (8.6)
Straightforwardly, we obtain the decay rate as
$\Gamma(\phi\rightarrow\bar{\psi}\psi)\simeq\frac{3}{4\pi}\frac{m_{\psi}^{2}M}{M_{p}^{2}},$ (8.7)
and the reheating temperature as
$T_{\text{reh}}\simeq 8.07\times 10^{-4}m_{\psi}.$ (8.8)
By using the ACT reheating temperature constraint of $10^{4}$ GeV, $m_{\psi}\sim 10^{8}$ GeV could satisfy the prediction.
8.3 Remarks on the perturbative mode
In our discussion about the perturbative reheating by scalar $\chi$ and fermion $\psi$, it is shown that the scalar mode is disfavored by the ACT. Furthermore, as it is enhanced by BEC, the reheating temperature from this mode could be larger than $\sim 10^{9}$ GeV. On the other hand, the reheating mode via fermions is more favorable, as it can decay to SM particles due to their chiral nature [42].
In Ref. [43], it is shown that both reheating mechanisms are rather oversimplified, as they neglect the violation of the fluctuation-dissipation theorem (see Refs. [44, 45, 46]). Additionally, these mechanisms overlook the effects of preheating and can only be significant if preheating is inefficient. This regime slowly transitions to a radiation-dominated regime, and the perturbative decay channel opens only gradually. It is implied that perturbative reheating via bosons is strongly disfavored due to the high reheating temperature, which contradicts the low reheating temperature prediction. Also, during the matter-dominated regime, all dominant particle production is mostly heavy non-relativistic [40]. On the other hand, fermionic reheating suffers not only from the same problem but also from the Pauli exclusion principle, which further suppresses their production. Even though this mode is more favorable than the bosonic one, its reheating scenario is still not convincing.
In addition, Ref. [30], which discusses inflation with a non-minimal coupling model, shows that perturbative reheating is strongly disfavored. This is because achieving a favorable reheating temperature requires a violation of unitarity (i.e., $\xi\gg 1$).
8.4 Reheating temperature by preheating
As we learn from the perturbative reheating, it has some issues at certain points. Thus, the reheating temperature should be obtained from the burst of particle production due to resonance. However, before we proceed, we need to assume the most possible mechanism.
Our chosen mechanism is as follows: Due to the nature of the matter-dominated regime, it is expected that the reheating temperature will be evaluated at the end of the preheating stage. However, the reheating temperature should not be evaluated from the ordinary perturbative decay of either bosons or fermions, as discussed in the previous subsection. Instead, the reheating temperature should be evaluated from the last burst of scalaron oscillation that produces scalar $\chi$ at the end of the preheating stage [^9] This was discussed in our previous paper (see Ref. [30]), although we used the beginning of the second stage of preheating.. The produced scalar $\chi$ will immediately decay into relativistic fermions $\psi$ to reheat the universe [^10] We choose the bosons $\chi$ as the product at the zero crossing of the scalaron due to the abundance of bosons over fermions during preheating. Fermions $\psi$ are chosen as the decay product of $\chi$ because they are more directly related to SM particles due to their chiral nature.. Intuitively, we can assume that during this time, the remaining energy density of the inflaton is very small. Thus, a low reheating temperature is expected.
To calculate this reheating temperature, it is written as
$T_{\text{reh}}=\left(\frac{30}{g^{*}\pi^{2}}\delta\rho\right)^{1/4},$ (8.9)
where $\delta\rho$ is the energy density produced by the single crossing during the ’last’ preheating. The $\delta\rho$ can be evaluated as [30, 31]
[…]
where
$n_{\chi}=\int\frac{d^{3}k_{\chi}}{(2\pi)^{3}}e^{2\mu_{p}Mt}\simeq\frac{k_{\chi}^{3}}{6\pi^{2}}$ (8.11)
is the number density of $\chi$. We set $\mu_{p}\simeq 0$ since during this time the preheating is almost ended. The mass $\bar{m}_{\chi}$ is estimated from the Eq. (3.5) as
$|\bar{m}_{\chi}|=\sqrt{\sqrt{6}\frac{\xi M^{2}\tilde{\phi}}{M_{p}}\sin(Mt)}.$ (8.12)
To complete Eq. (8.10) we should consider that there is an interaction between $\chi$ and $\psi$ via
$-y\chi\bar{\psi}\psi,$ (8.13)
where $y$ is the Yukawa coupling between $\chi$ and $\bar{\psi}\psi$. The decay rate for this channel is approximately given by,
$\Gamma_{\chi\rightarrow\bar{\psi}\psi}\simeq\frac{y^{2}m_{\chi}}{8\pi}.$ (8.14)
With these, we can solve the Eq. (8.10) as
$\delta\rho\simeq 3.71\times 10^{-6}y^{2}k_{\chi}^{3}m_{\chi}\sqrt{\xi}$ (8.15)
Straightforwardly, the reheating temperature can be estimated as
$T_{\text{reh}}\simeq 1.83\times 10^{-2}\left(y^{2}k_{\chi}^{3}m_{\chi}\sqrt{\xi}\right)^{1/4}\quad\text{GeV}.$ (8.16)
The reheating temperature is now shown to be dependent on several parameters. In short, one can see the Table. 2. It is shown that low $y$, $k_{\chi}$, and $m_{\chi}$ are required to achieve the favored reheating temperature of $\sim 10^{4}$ GeV.
$y$ $k_{\chi}$ $m_{\chi}$ $T_{\text{reh}}$
$10^{-1}$ $10^{-9}M_{p}$ $10^{-16}M_{p}$ $4.46\times 10^{5}$ GeV
$10^{-2}$ $10^{-9}M_{p}$ $10^{-16}M_{p}$ $1.41\times 10^{5}$ GeV
$10^{-3}$ $10^{-10}M_{p}$ $10^{-12}M_{p}$ $4.46\times 10^{4}$ GeV
$10^{-4}$ $10^{-10}M_{p}$ $10^{-12}M_{p}$ $1.41\times 10^{4}$ GeV
$10^{-4}$ $10^{-9.5}M_{p}$ $10^{-12}M_{p}$ $3.34\times 10^{4}$ GeV
$10^{-4}$ $10^{-9}M_{p}$ $10^{-12}M_{p}$ $7.92\times 10^{4}$ GeV
$10^{-4}$ $10^{-8.5}M_{p}$ $10^{-12}M_{p}$ $1.88\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-12}M_{p}$ $4.46\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-13}M_{p}$ $2.51\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-14}M_{p}$ $1.4\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-15}M_{p}$ $7.92\times 10^{4}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-16}M_{p}$ $4.46\times 10^{4}$ GeV
Table 2: The reheating temperature with varying $y$, $k_{\chi}$ and $m_{\chi}$. We set $\xi=1$ and adjust the parameters to be in the range of $T_{\text{reh}}$ of $10^{4}\sim 10^{5}$ GeV.
Considering the chosen parameters, we need to clarify some favored values. The coupling $y$ should be small to suppress instant reheating (see [19]) [^11] Fortunately, if we set $y$ to be small, our coupled e.o.m. (3.2) and (3.3) do not necessarily consider the growth of $\psi$.. For $m_{\chi}=10^{-15}M_{p}\sim 10^{-16}M_{p}$, this lies close to the electroweak scale. A model with this mass range favors linkage to the SM. Interestingly, a mass of this value does not contradict our model. In addition, since $m_{\psi}\ll m_{\chi}$ is required to allow for smooth decay, it is clear that $m_{\psi}$ is on the order of SM particle masses. For our chosen parameters, $\chi$ could correspond to the Higgs boson, and $\psi$ could be SM fermions [^12] The small Yukawa coupling may naturally disfavor some large quark couplings.. However, we must be aware that due to the large couplings of the Higgs to gauge bosons (and some large-coupling fermions), such interactions could spoil our low-reheating inflationary model. Finally, it should be noted that both $\chi$ and $\psi$ masses should lie above the electroweak scale, as neither fields do not behave like an SM particle. Note that $y$ could be much smaller. Thus, all other parameters could be set to higher values, $k_{\chi}\gg 10^{-8}M_{p}$ and $m_{\chi}\gg 10^{-16}M_{p}$.
9 Conclusion
In this paper, we reinvestigate Starobinsky inflation, particularly the preheating stage, following the release of the ACT results. These results suggest a significant increase in the e-fold number ($N\simeq 75$) and constrain the reheating temperature to $T_{\text{reh}}\sim 10^{4}$ GeV if we adopt the $68\%$ confidence level. Consequently, the preheating and reheating mechanisms must be modified to align with these new results. Such a low reheating temperature affects the entire preheating stage, not just the reheating mechanism. In addition, the direct perturbative decay from the scalaron to scalar $\chi$ is strongly disfavored if the preheating mechanism is dominant and the high reheating temperature is implied. On the other hand, reheating via the decay of the scalaron to fermions $\psi$ is also disfavored if the preheating stage is dominant, although less severely, due to the possibility of achieving a lower reheating temperature and its chirality nature.
Ultimately, the reheating mechanism is realized through a specific preheating process. During resonance, the oscillating scalaron produces the scalar $\chi$. While fermion production is also possible, it is heavily suppressed due to the Pauli exclusion principle. During the preheating stage, perturbative decay of the scalaron and $\chi$ occurs but proceeds very slowly and does not contribute adequately to the reheating temperature. At the end of the preheating stage, the final burst of particle production from the scalaron to $\chi$, followed by the immediate decay of $\chi$ into fermions $\psi$, becomes the sole contributor to the reheating temperature. The reheating mechanism via this mode is found to be consistent with our chosen parameters.
In the preheating stage, we consider two mechanisms, which we refer to as Case 1 and Case 2, distinguished by the presence of tachyonic conditions. Case 1 describes the preheating stage when the tachyonic condition is not applied. In this mode, preheating is shown to be inefficient. Consequently, the preheating stage lasts much longer than the analytical prediction ($\gg 10^{5}\pi/M$). On the contrary, Case 2 exhibits purely tachyonic conditions. In this mode, the preheating stage ceases rapidly due to efficient particle production ($\ll 10^{5}\pi/M$). Neither the numerical prediction from Case 1 nor that from Case 2 alone satisfies the analytical calculation. For this reason, the realistic preheating mechanism may involve a combination of both Case 1 and Case 2. To accommodate this idea, at least one momentum, $k_{\phi}$ or $k_{\chi}$, should not be fixed but instead follow a time-dependent function that disrupts the tachyonic conditions. This is referred to as the ceased-out mechanism [30].
The disruption of the tachyonic conditions can occur through two mechanisms, based on the following question: At which stage does tachyonic preheating occur—during the early or late phase of preheating? This question arises particularly when we consider PBH production during preheating. It has been shown that the appropriate PBH production occurs during the early stage of preheating. Based on this idea, it is clear that a preheating stage where tachyonic condition intervenes on the late stage is strongly disfavored.
We have also considered the favored parameters in our model based on our discussion in the previous sections. In our model, $y$ should be small to avoid instant reheating, which would lead to a high reheating temperature. The non-minimal coupling $\xi$ should be of order unity or much lower due to unitarity issues (in this paper, we set $\xi=1$). The mass of $\chi$ should be lower than $\frac{1}{2}M$ to ensure that the scalaron decays completely. In addition, $m_{\chi}$ should be higher than the electroweak scale ($\gtrsim 10^{3}$ GeV $\approx 10^{-15}M_{p}$), as $\chi$ is shown to be different to the Higgs field to preserve the viability of our model [^13] If $m_{\chi}$ is at the electroweak scale and not the Higgs field, it may have already been observed.. The mass $m_{\psi}$ should be much lower than $m_{\chi}$ to ensure a successful reheating scenario [^14] As $\psi$ is supposed to reheat the universe, it should be relativistic. Thus, $m_{\psi}\ll m_{\chi}$ is preferred.. The parameters $k_{\phi}$ and $k_{\chi}$ in the early stage of preheating should satisfy the tachyonic condition (see eq. (3.12)). One should note that, as $\chi$ in the early stage of preheating is heavy non-relativistic, it is implied that $k_{\chi}$ should be low. On the contrary, in the late stage, they should violate the tachyonic condition. Finally, the evolution of $k_{\phi}$ is favorable for PBH production. One can see table 3 for the appropriate values of the parameters. Note that all parameters in 3 are under the constraint of the reheating temperature.
Parameter value Note
$y$ $\ll 10^{-3}$ avoid instant preheating
$m_{\chi}$ $10^{-6}\sim 10^{-15}M_{p}$ $<\frac{1}{2}M$
$m_{\psi}$ $10^{-7}\sim 10^{-15}M_{p}$ $\ll m_{\chi}$
$k_{\phi}k_{\chi}$ $>8.23\times 10^{-11}M_{p}^{2}$ $>\xi M^{2}$, early preheating
$k_{\phi}k_{\chi}$ $<8.23\times 10^{-11}M_{p}^{2}$ $<\xi M^{2}$, end preheating
Table 3: The viable values of the parameters in our model. Note that we set $\xi=1$.