### v3

# The Preheating Stage on The Starobinsky Inflation after ACT

Abstract
In this paper, we reinvestigate the Starobinsky inflation model and its reheating features in light of the recent ACT results. To make the Starobinsky model consistent with the ACT data at the $68\%$ confidence level, the number of e-folds must increase while the reheating temperature decreases. We find that the Starobinsky model requires a spectator field to achieve efficient preheating. The preheating stage and the reheating temperature must be significantly adjusted to accommodate the lower temperature. In this paper, the favored non-minimal coupling of the produced particles is approximately $10$ or slightly lower. We also present viable parameter sets that fit the preferred reheating mechanism in this model. For certain parameter choices, the daughter fields could potentially be detected in future collider experiments such as the LHC or the ILC. Furthermore, our proposed mechanism can reproduce the lower reheating temperature, but it fails when the temperature falls below $1$ GeV.
1 Introduction
The Starobinsky model [starobinsky], as the simplest example of a modified gravity theory [de2010f, sotiriou2010f, nojiri2011unified], is considered one of the most successful inflationary models. Its simplicity lies in the fact that it relies solely on gravity as the driving mechanism for inflation. Despite this simplicity, it remains consistent with the Cosmic Microwave Background (CMB) observations from Planck data [cmb]. In recent years, this model and its various extensions have continued to attract considerable attention and discussion [kehagias, Toyama:2024ugg, Ketov:2024klm, Ketov:2025nkr, He:2024wqv, He:2018gyf, Risdianto:2022ium, Sebastiani:2013eqa, Bamba:2014jia, Nojiri:2017ncd, CruzDombriz:2016wcb, Odintsov:2018zhw, Odintsov:2022hxu, Odintsov:2021kup, Odintsov:2020thl, Nojiri:2019fft, Elizalde:2018rmz]. However, the Starobinsky model is now challenged by newly established observational data from the Atacama Cosmology Telescope (ACT) [louis2025atacama]. This new data shows that the slow-roll parameter is $n_{s}=0.9709\pm 0.0038$. When combined with measurements from CMB lensing, Baryon Acoustic Oscillation (BAO), and the Dark Energy Spectroscopic Instrument (DESI), the extended result yields $n_{s}=0.9743\pm 0.0034$. This result indicates that the Starobinsky model is excluded at the $2\sigma$ level for $N\approx 60$. The model remains viable if a large number of e-folds ($N\approx 75$) is allowed, which leads to a lower reheating temperature [Drees:2025ngb, liu2025reconciling, mohammadi2025starobinsky, zharov2025reheating, gialamas2025keeping]. Alternatively, this issue can be addressed by introducing higher-curvature perturbations to satisfy the new constraint [addazi2025curvature].
The reheating temperature of the Starobinsky model is found to be relatively low $\sim 10^{9}$ GeV [de2010f, dorsch2024gravitational] compared to other models, such as Higgs inflation [bezrukov2009initial]. This lower reheating temperature is often preferred in inflationary cosmology, as it helps avoid the overproduction of unwanted relics [kallosh2000gravitino]. For example, Refs. [ellis1982inflation, kawasaki1995gravitino] argue [^1] See also Ref. [mcdonald2000reheating]. that the reheating temperature should be below $10^{9}$ GeV to prevent excessive production of gravitino. Such a low reheating temperature significantly influences the thermal history of the universe, particularly with respect to Baryogenesis [davidson2000baryogenesis] and the production of dark matter. However, as for leptogenesis, it requires at least several MeV of reheating temperatures to operate successfully [hannestad2004lowest, choi2017new]. In summary, estimation of the reheating temperature range is wide, from a few MeV up to nearly the Planck scale. This broad range suggests that the reheating temperature is still poorly constrained. In addition to the ACT observation results, the Starobinsky inflation model is expected to yield a much lower reheating temperature.
The preheating stage of the Starobinsky model has received comparatively less attention in the literature than its inflationary phase. This is largely because the inflationary dynamics are better constrained, whereas the preheating phase tends to be more model-dependent [van2017reheating, fu2019nonlinear]. [^2] One can see the historical preheating stage in Refs. [Dolgov:1989us, Traschen:1990sw, Shtanov:1994ce, Kofman:1994rk] In the Starobinsky model, preheating is primarily driven by the interaction of the produced particles with gravity. While this makes it a simple and potentially promising scenario, our limited understanding of gravity makes it more challenging to analyze fully. In this paper, we investigate the preheating mechanism in the Starobinsky model, focusing on the efficient energy drain by preheating, the constraints from primordial black hole (PBH) production, and the resulting reheating temperature.
This paper is organized as follows: In Section 2, we discuss the inflationary description of the Starobinsky model. In Section 3, we estimate the resulting reheating temperature by the ACT constraint. In Section 4, we examine the preheating phase of the Starobinsky model, which is non-minimally coupled to scalars. In Section 5, we consider the possibility of fermionic preheating in our model. We also discuss primordial black hole (PBH) formation in Section 6. The reheating scenario is discussed in Section 7. Finally, we present our conclusions in Section 8.
2 The inflationary model
We start this section by writing the action of the model in the Jordan frame as [^3] In this paper, we used (,+,+,+)(-,+,+,+) convention.
$S_{J}=\int d^{4}x\sqrt{-g_{J}}\left[\frac{1}{2}M_{p}^{2}f(R_{J})\right],\hskip 28.45274ptf(R_{J})=R_{J}+\frac{R^{2}_{J}}{6M^{2}},$ (2.1)
where $g_{J}$ is the determinant of the metric tensor $g^{\mu\nu}_{J}$, $M_{p}=1/\sqrt{8\pi G}$ is the reduced Planck mass, $R_{J}$ is the Ricci scalar, and $M$ is the effective inflaton’s mass at the end of inflation. The subscript $J$ represents the Jordan frame. By using the conformal transformation as
$g^{\mu\nu}_{E}\equiv\Omega^{-2}g^{\mu\nu}_{J},\hskip 28.45274pt\Omega^{2}\equiv\frac{df(R_{J})}{dR_{J}}=1+\frac{R_{J}}{3M^{2}}\equiv\exp\left({\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right),$ (2.2)
we obtain the action in the Einstein frame as follows
$S_{E}=\int d^{4}x\sqrt{-g_{E}}\left[\frac{1}{2}M_{p}^{2}R_{E}-\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)\right],$ (2.3)
where ${R}_{E}=R_{J}\Omega^{-2}-6\Omega^{-3}\square\Omega$ and
$V(\phi)=\frac{3}{4}M_{p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)^{2}.$ (2.4)
For the rest of this paper, we will omit the subscript $E$ for simplicity.
Firstly, we will write the scalar spectral index ($n_{s}$), tensor-to-scalar ratio ($r$), slow-roll parameters ($\epsilon$ and $\eta$) and the scalar power spectrum ($A_{s}$) respectively as
$n_{s}=1-6\epsilon+2\eta,\quad r=16\epsilon,\quad A_{s}=\frac{V(\phi)}{24\pi^{2}\epsilon}\quad\epsilon=\frac{1}{2}M_{p}^{2}\left(\frac{V^{\prime}}{V}\right)^{2}\quad\eta=M_{p}^{2}\frac{V^{\prime\prime}}{V}.$ (2.5)
The prime denotes the derivative with respect to scalaron $\phi$. By using the ACT from the extended dataset P-ACT-LB yields [louis2025atacama] $n_{s}=0.9743\pm 0.0038$, $A_{s}\approx 2.13\times 10^{-9}$ at the pivot scale $\kappa_{*}=0.05$ Mpc-1. It can be used to determine the new constraint on the $\phi_{\text{ini}}$, which corresponds to the field value of the scalaron at $\kappa_{*}$, to be
$\phi_{\text{ini}}\approx 5.72\hskip 2.84526ptM_{p},$ (2.6)
where this value is set by the center value of $n_{s}$, and the details may be explained shortly. The inflaton’s field value at the end of inflation is obtained by setting $\epsilon=1$, which corresponds to the violation of the slow-roll conditions as
$\phi_{\text{end}}\simeq 0.94\hskip 2.84526ptM_{p}.$ (2.7)
By using the last two results, we can obtain the number of e-folds $N$ as
$N=\frac{1}{M_{p}^{2}}\int^{\phi_{\text{ini}}}_{\phi_{\text{end}}}\frac{V}{V^{\prime}}\simeq 75.51,$ (2.8)
much higher than the previously predicted  [^4] One should also note that slight changes in the constraint of $n_{s}$ could potentially alter the results of other parameters greatly. and this large number of e-folds is expected. [^5] The chosen e-fold greater than 60 is rather implausible but still acceptable [liddle2003long]. Note that we only used the center value of $n_{s}$. By using the range of $n_{s}$, we will obtain an e-fold range of 668766-87, which will later significantly alter the reheating temperature. However, to simplify our analysis, we will only consider the center value of $n_{s}$.This assumption is made to incorporate the new mechanism consistent with the lower reheating temperature indicated by the recent ACT results.
In addition, the effective mass of the scalaron $M$ is also affected by the new ACT constraint on the scalar power spectrum $A_{s}$ as
$M\simeq 9.07\times 10^{-6}M_{p},$ (2.9)
which is slightly lower than if constrained by the Planck result [cmb].
3 The reheating constraint by ACT results
In this, we will estimate the reheating temperature with the new ACT results. This constrained reheating temperature will be important to estimate the proper reheating process as depicted in section 3. Straightforwardly, it is necessary to write the formula as
$\frac{\kappa_{*}}{a_{0}H_{0}}=\frac{a_{\kappa}}{a_{\text{end}}}\frac{a_{\text{end}}}{a_{\text{reh}}}\frac{a_{\text{reh}}}{a_{\text{eq}}}\frac{a_{\text{eq}}}{a_{0}}\frac{H_{\text{eq}}}{H_{0}}\frac{H_{\kappa}}{H_{\text{eq}}},$ (3.1)
where $a_{0}=1$ and $H_{0}=66.78\hskip 2.84526pt\text{km Mpc}^{-1}\text{s}^{-1}$ [louis2025atacama] represent the scale factor and Hubble parameter today. $a_{\kappa}$ and $H_{\kappa}$ represent the scale factor and Hubble parameter at the pivot scale $\kappa_{*}=0.05\hskip 2.84526pt\text{Mpc}^{-1}$. The subscripts ’eq’, ’end’, and ’reh’ correspond to ’matter-radiation equality’, ’end of inflation’, and ’reheating’, respectively. Note that we solve Eq. (3.1) by using $T_{\text{eq}}=0.8$ eV, and the end of preheating is evaluated at $\tilde{\phi}=M$. Finally, by using (3.1) with e-fold $N=75.5$, we obtain an upper bound for the reheating temperature as $T_{\text{reh}}\sim 10^{4}$ GeV. This value is five orders of magnitude lower than the previous prediction [de2010f]. The reheating temperature in our calculation is evaluated at the early radiation-dominated regime.
Note that we used the center value of $n_{s}$. This is to simplify our preheating mechanism. The small change in $n_{s}$ will greatly change the e-fold and finally alter the reheating temperature more severely. The reheating temperature should be directly connected to the preheating stage mechanisms. In that case,
4 The preheating
When $\phi\lesssim M_{P}$, the inflaton oscillates, marking the end of inflation. The particle produced by this mode is non-perturbative. If we assume $\chi$ is dominantly produced during this mode, which is presented by the action:
$S_{\chi}=\int d^{4}x\sqrt{-g_{J}}\left[\frac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\frac{1}{2}m_{\chi}^{2}\chi^{2}-\xi R_{J}\chi^{2}\right],$ (4.1)
we can evaluate its particle production. Firstly, we will calculate the analytical solution of $\phi$ as [kofman1997towards]
Figure 1: The plot of $\phi$ with time $t$. The unit of $\phi$ is $M_{p}$ while the unit of $t$ is in $M_{p}^{-1}$.
$\phi(t)=\tilde{\phi}(t)\sin(Mt),\hskip 28.45274pt\tilde{\phi}(t)=2\sqrt{\frac{2}{3}}\frac{M_{p}}{Mt}.$ (4.2)
On the other hand, the equation of motion (e.o.m.) of $\chi$ can be written by varying $S+S_{\chi}$ by $\delta\chi$ as
$\ddot{\chi}_{k}+3H\dot{\chi}_{k}-3\sqrt{\frac{2}{3}}\frac{\dot{\phi}\dot{\chi}_{k}}{M_{p}}+\left[\frac{k^{2}_{\chi}}{a^{2}}+e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}m_{\chi}^{2}-3\xi M^{2}e^{2\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}-1\right)\right]\chi_{k}=0,$ (4.3)
where we used the Heisenberg representation as
$\chi(\textbf{x},t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}\left(\hat{a}_{k}\chi_{k}(t)e^{-i\textbf{k}_{\chi}\cdot\textbf{x}}+\hat{a}^{\dagger}_{k}\chi^{*}_{k}(t)e^{i\textbf{k}_{\chi}\cdot\textbf{x}}\right),$ (4.4)
on calculating (4.3). In Fig. 2, we show that the resonance production of $\chi_{k}$ is highly sensitive to both the non-minimal coupling $\xi$ and the initial condition $\chi_{k}(0)$. Our results indicate that particle production from zero-crossings alone is inefficient. Instead, efficient preheating relies primarily on tachyonic instability immediately after the first crossing. As we stated earlier, tachyonic preheating in the Starobinsky model is also strongly dependent on the initial conditions set at the start of preheating. If these conditions are not satisfied, Starobinsky inflation faces two possible results: (1) inconsistency with the observed particle abundance today (especially if $\chi$ is related to dark matter) or (2) prolonged preheating to generate the required number of particles.
Figure 2: The growth of $\chi_{k}$ by the resonance production. We fixed $M=9.07\times 10^{-6}M_{p}$, $M_{p}=1$, $k_{\chi}=10^{-8}M_{p}$, $m_{\chi}=10^{-8}M_{p}$, and $a=1$. The solid-red represents to the $\chi_{k}(0)=0$, The dashed-blue represents to the $\chi_{k}(0)=10^{-5}M_{p}$, and The solid-black represents to the $\chi_{k}(0)=10^{-2.5}M_{p}$. The varied $\xi$ is depicted in the corresponding panels: panel (A) uses $\xi=1$, panel (B) uses $\xi=4$, panel (C) uses $\xi=7$, and panel (D) uses $\xi=10$. The $\chi_{k}$ is in a unit of $M_{p}$ while $t$ is in a unit of $M_{p}^{-1}$.
The initial conditions can be provided via four mechanisms: perturbative decay at an early stage, scattering of the scalaron, quantum fluctuation $\delta\chi$, and $\chi$ as a spectator field during inflation. Among the four mechanisms, we will qualitatively analyze which mechanism is the most favored.
At the early stage of preheating, the inflaton perturbative decay and the scattering process are possible. The single perturbative decay, in principle, is more favored than scattering, as it needs a large density of incoherent inflaton particles to scatter. In contrast, at the start of preheating, the inflaton is mostly a coherent oscillating field. The perturbative decay itself is slow to be produced at the early stage of preheating. Additionally, if the decay occurs during inflation, it may spoil the slow-roll condition or leave observational relics. On the other hand, the favored $\chi\sim 10^{-2.5}M_{p}$ could be obtained from the fluctuation of $\delta\chi$. But it can only provide about $H/2\pi\sim 10^{-6}M_{p}$, in which about $10^{-3}$ of smaller order from the desired value. Lastly, the only possible way came from the fact that $\chi$ should be the spectator field during inflation. On the other hand, as $\chi$ is small-coupled with scalaron, it can be assumed it is safe as a spectator field. We can also conclude that for the successful preheating of the Starobinsky inflation, the appearance of the spectator field is necessary. At the early stage of preheating, in order to fulfill the non-relativistic conditions and the efficient preheating, $k_{\chi}$ should be extremely small. Thus, it gives $k_{\chi}^{2}\ll\xi M^{2}$.
Before proceeding, we clarify why the analytical time evolution of $\phi$ is used, even though it could in principle be numerically coupled with $\chi$ to obtain exact results. First, in the pure Starobinsky model, the scalaron $\phi$ is only weakly coupled to other fields during the initial stage of preheating. In the Einstein frame, the Lagrangian of the $\chi$ field is heavily suppressed by the factor $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}$, which indicates that $\phi$ and $\chi$ are weakly coupled. This justifies treating $\chi$ as a spectator field during inflation, so that the analytical form of Eq. (4.2) remains valid until $\phi$ becomes small. Also, we restrict our attention to the early phase, where $\phi$ is still not too small after the first crossing. In this regime, the analytical estimation of the time evolution of $\phi$ remains reliable. Second, solving the coupled equations of motion poses a significant challenge, as it involves complicated calculations, including integro-differential equations in $\chi^{2}$. Third, since our paper primarily focuses on the reheating mechanism in light of the new ACT results, we approximate preheating by considering how efficiently it can drain the inflaton field energy density. A more detailed analysis of the preheating stage may be presented in future work.
In the following, we will describe the analytical approximation to dig further into the particle production of $\chi_{k}$. At the end of inflation, Eq. (4.3) can be approximated by using $a=1$ and evaluated when $\tilde{\phi}<M_{p}$ as
$\ddot{\chi}_{k}+\Bigg[\left(k^{2}_{\chi}+m_{\chi}^{2}\right)+\left(\sqrt{\frac{2}{3}}\frac{m_{\chi}^{2}+3\xi M^{2}}{M_{p}}\right)\tilde{\phi}\sin(Mt)\Bigg]\chi_{k}\simeq 0.$ (4.5)
The approximation given in Eq. (4.5) can be considered to be oversimplified. As shown in the Fig. 3, the analytical results are heavily deviated from Eq. (4.3). In that figure, we used the comparison of $\chi_{k}^{2}$ instead of $\chi_{k}$. In Fig. 3, we used the amplification factor of $\chi_{k}$ as $\sim 10$ (or $\sim 10^{2}$ for $\chi^{2}_{k}$) on the numerical result to be aligned with the analytical result. We will use this information to estimate the particle production during the crossing and the deviation factor of the Mathieu instability.
Figure 3: The comparison of analytical result (solid-red) and numerical result (solid-black) in $\chi_{k}^{2}$. We used amplification $1.1\times 10^{1.8}$ on the numerical result to match the analytical ones. In this plot, we used $\xi=10$, $k_{\chi}=10^{-8}M_{p}$, $m_{\chi}=10^{-8}M_{p}$, $\chi_{k}(0)=10^{2.5}M_{p}$ and $\dot{\chi}_{k}=0$.
For the near-zero crossing ($\sin(Mt)\simeq Mt$), we can approximate Eq. (4.5) as
$\frac{d^{2}{\chi}_{k}}{dq^{2}}+\left(p^{2}+q\right)\chi_{k}=0,$ (4.6)
where
$\begin{split}&p^{2}=\left(k_{\chi}^{2}+m_{\chi}^{2}\right)\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{-2/3}\\ &q=\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{1/3}t.\end{split}$ (4.7)
This simplification will help us to understand the numerical results depicted in Fig. 2, which is obtained by solving Eq. (4.3).
For $p^{2}\ll q$, we can estimate the particle production of $\chi$ by the first crossing of $\phi$ as [bezrukov2009initial, risdianto2025second]
$\delta\rho=\int^{\infty}_{0}\frac{d^{3}k_{\chi}}{(2\pi)^{3}}e^{-\pi p^{2}}\bar{m}$ (4.8)
where
$\begin{split}&\bar{m}\simeq\sqrt{\sqrt{6}\frac{\xi M^{2}\tilde{\phi}}{M_{P}}}\end{split}.$ (4.9)
Analytically, by calculating Eq. (4.8), we obtain that in the first zero crossing (after the end of inflation), it produces
$\delta\rho_{\text{cross}}\simeq 0.26\times\xi^{3/2}M^{4}\sim 10^{-21}\xi^{3/2}M_{p}^{4}.$ (4.10)
In this calculation, for the crossing to achieve the same energy as the inflaton ($\tfrac{1}{2}M^{2}\tilde{\phi}^{2}\sim 10^{-11}M_{p}^{4}$), we require $\xi\gg\mathcal{O}(1)$, which is strongly disfavored due to naturalness issues. In addition, Fig. 3 shows that the actual value of $\delta\rho_{\text{cross}}$ obtained analytically is about $\sim 10^{2}$ times smaller, indicating that the inflaton’s energy drain through zero-crossing becomes less efficient. It should also be noted that $\delta\rho_{\text{cross}}\propto|\chi_{k}|^{2}$.
Lastly, we will describe the e.o.m. of $\chi_{k}$ in the redefined analytical result. Straightforwardly, Eq. (4.5) can be shown by
$\begin{split}&Mt\equiv 2z-\pi/2,\hskip 28.45274ptA\equiv 4\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{M^{2}}\right),\\ &Q\equiv\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}.\end{split}$ (4.11)
So we obtain the famous Mathieu equation as
$\frac{d^{2}\chi_{k}}{dz^{2}}+(A-2Q\cos(2z))\chi_{k}=0.$ (4.12)
The numerical calculation of the Mathieu instability is shown in Fig. 4. The growth of the $\chi_{k}$ is depicted by the characteristic exponent $\mu_{p}$, which can be written analytically as
$\mu_{p}=\frac{1}{2\pi}\ln\left(1+2e^{-\pi p^{2}}-2\sin\theta_{\mu}e^{-\frac{\pi p^{2}}{2}}\sqrt{1+2e^{-\pi p^{2}}}\right),$ (4.13)
where $\theta_{\mu}$ is a phase. It is shown in the Mathieu instability band of Fig. 4 that certain parameters could potentially fall in the instability region in which the particle is largely produced. However, this semi-analytical result of the instability band could not show the effect of the initial conditions. Also, the parameter used in the numerical results could fit in the Fig. 4 with much smaller amplification by an order of 10.
Figure 4: Mathieu instability region related to the semi-analytical result depicted in Eq. (4.12). The right legend corresponds to the value of $\mu_{p}$. The white region corresponds to the stable condition.
5 On the fermionic preheating
In this part, we will discuss the possibility that, instead of the scalar $\chi$, the scalaron is coupled with fermions. So far, we have only discussed the bosonic preheating, even though fermionic preheating is equally possible. This type of preheating has already been discussed in, e.g., Refs. [greene2000fermionic, greene1999fermion, giudice1999fermion, peloso2000fermion]. In this section, we will qualitatively discuss the possibility of this type of preheating.
If the scalaron is only coupled with certain fermions, at a large coupling [greene1999fermion], the preheating is similar to the bosonic case. However, in our paper, it is heavily suppressed by the Pauli exclusion principle, in contrast to the Bose-Einstein condensate (BEC) that can occur during bosonic preheating. Fermion production becomes less suppressed once the universe expands. However, even after the end of preheating, the potential of the scalaron remains quadratic ($\propto\phi^{2}$), which still favors a matter-dominated regime. Fermions are more likely to be produced at the end of the preheating phase due to the looseness of the Pauli blocking. Finally, we safely said that preheating via this mode is strongly disfavored.
6 The PBH formation from preheating
In this section, we will discuss PBH formation due to preheating in the Starobinsky model. We will show later that the PBH formation is consistent with it being produced during the early stage of preheating.
Straightforwardly, the constraint on the PBH can be written as
$\beta_{\text{PBH}}=\frac{\rho_{\text{pbh}}}{\rho_{\text{tot}}}=\int^{\infty}_{\delta_{c}}P(\delta)d\delta,$ (6.1)
where $P(\delta)$ represents the probability distribution function, and it can be written as
$P(\delta)=\frac{1}{\sqrt{2\pi}\sigma_{\delta}}\exp\left(-\frac{\delta^{2}}{2\sigma_{\delta}^{2}}\right).$ (6.2)
By using $\beta_{\text{PBH}}=10^{-20}$ and $\delta=\delta_{c}=0.7$ [green2001primordial, niemeyer1999dynamics] on Eq. (6.1), we obtain the mass variance $\sigma_{\delta}$ as
$\sigma_{\delta}\simeq 0.08.$ (6.3)
However, we need to skip this part before putting the constraint into our parameters.
For the production of PBH during preheating, the strong parametric resonance should be applied. Our assumptions implied that PBH is most likely produced when [liddle2000super, green2001primordial]
$|Q|\equiv\left|\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}\right|\gg A.$ (6.4)
The power spectrum of the variable $x$ can be expressed as [liddle2000cosmological]
$\mathcal{P}_{(x)}=\frac{k_{x}^{3}}{2\pi^{2}}\braket{|x_{k}|{}^{3}},$ (6.5)
where $k_{x}=|\textbf{k}_{x}|$ represents the comoving wavenumber associated with $x$. Note that $x_{k}$ denotes the inverse Fourier transform of $x$. The influence of preheating on $\chi$ production can be determined as [kofman1997towards, liddle2000super, green2001primordial]
$\mathcal{P}_{\delta\chi}=\mathcal{P}_{\delta\chi}|_{\text{end}}\exp\left(2\mu_{p}M\Delta t\right),$ (6.6)
where $\Delta t$ corresponds to the time required to obtain the appropriate number of PBH. The characteristic exponent $\mu_{p}$ here is evaluated at $\theta_{\mu}=0$, in which $\mu_{p}=(2\pi)^{-1}\ln(1+2\exp(-\pi p^{2}))$, where $p$ is taken from Eq. (4.7) as
$p^{2}=\frac{k_{\chi}^{2}+m_{\chi}^{2}}{\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{2/3}}=\frac{1}{18\sqrt{Q}}\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{k_{\text{end}}^{2}}\right),$ (6.7)
where we introduce
$k_{\text{end}}^{2}=\frac{1}{18\sqrt{2}}\left|\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}\right]^{1/6}\right|M^{5/3}$ (6.8)
as the comoving wavenumber evaluated at the Hubble exit at the end of inflation. For $Q\gg 1$, in which it is required to get the largest production, we get $\mu_{p}=\mu_{0}\simeq\frac{1}{2\pi}\ln 3$. In addition, we will calculate the non-adiabatic power spectrum of $\chi$ as [liddle2000super, green2001primordial]
$\mathcal{P}_{\zeta_{\text{n-ad}}}(k_{\chi})\simeq\mathcal{A}\left(\frac{k_{\chi}}{k_{\text{end}}}\right)^{3}I(p,M\Delta t),$ (6.9)
where [^6] The third bracket is coming from the interaction coupling between $\phi$ and $\chi$ via $\mathcal{L}_{\text{int}}=\sqrt{\frac{2}{3}}\frac{m^{2}_{\chi}}{M_{P}}\phi\chi^{2}$ divided by mass scale $M$.
$\mathcal{A}\equiv\frac{3\times 2^{3/2}}{\pi^{6}\mu_{p}^{2}}\left(\frac{\tilde{\phi}}{M_{p}}\right)^{2}\left(\frac{H_{\text{end}}}{M}\right)^{4}\left(\sqrt{\frac{2}{3}}\frac{m^{2}_{\chi}}{MM_{P}}\right)^{2}Q^{-1/4}$ (6.10)
and
$I(p,M\Delta t)\equiv\frac{3}{2}\int^{\infty}_{0}p^{\prime 2}dp^{\prime}\int^{\pi}_{0}\sin\vartheta d\vartheta e^{2(\mu_{p^{\prime}}+\mu_{p-p^{\prime}})M\Delta t}.$ (6.11)
Note that $\vartheta$ is the angle between $p$ and $p^{\prime}$. To solve the last equation, we need to impose [green2001primordial]
$\mu_{p^{\prime}}\simeq\mu_{p-p^{\prime}}\approx\mu_{0}+\frac{p^{\prime 2}}{3}.$ (6.12)
So we obtain
[…]
Finally, the mass variance $\sigma_{\delta}$ during the matter-dominated era can be calculated as follows [green2004NewCalculation]
$\sigma_{\delta}^{2}=\frac{16}{81}\int^{\infty}_{0}W^{2}(\tilde{k}_{\chi},\mathcal{R})\left(\frac{\tilde{k}_{\chi}}{{k}_{\chi}}\right)^{4}\mathcal{P}_{\zeta_{\text{n-ad}}}\frac{d\tilde{k}_{\chi}}{\tilde{k}_{\chi}},$ (6.14)
where we introduced the window function $W(\tilde{k}_{\chi},\mathcal{R})=\exp\left(-\tilde{k}_{\chi}^{2}\mathcal{R}^{2}/2\right)$, with $\mathcal{R}=1/k_{\psi}$. Finally, we can use $\sigma_{\delta}$ calculated from Eqs.(6.14), (6.1) and (6.3) to obtain the constraint on PBH production to be
$0.0755^{2}\gtrsim\frac{0.02\cdot e^{0.6992\cdot M\Delta t}}{(M\Delta t)^{3/2}}\cdot\frac{m^{4}_{\chi}k_{\chi}^{3}}{\sqrt{\xi}M^{5}M_{p}^{2}}$ (6.15)
$m_{\chi}$ $k_{\chi}$ $\Delta t$ ($M_{p}^{-1}$)
$10^{-12}M_{p}$ $10^{-12}M_{p}$ $2.24\times 10^{7}$
$^{-12}M_{p} 1010$$^{-10}M_{p}$ $2.02\times 10^{7}$
$10^{-12}M_{p}$ $10^{-8}M_{p}$ $1.80\times 10^{7}$
$10^{-8}M_{p}$ $10^{-12}M_{p}$ $1.65\times 10^{7}$
$^{-8}M_{p} 1010$$^{-10}M_{p}$ $1.43\times 10^{7}$
$10^{-8}M_{p}$ $10^{-8}M_{p}$ $1.21\times 10^{7}$
$10^{-6}M_{p}$ $10^{-12}M_{p}$ $1.36\times 10^{7}$
$^{-6}M_{p} 1010$$^{-10}M_{p}$ $1.136\times 10^{7}$
$10^{-6}M_{p}$ $10^{-8}M_{p}$ $9.12\times 10^{6}$
Table 1: The duration of the required PBH abundance. In this table, we used $\xi=1$. Based on the table, it seems that the $\Delta t$ is insensitive to the chosen parameters.
Eq. (6.15) shows the dependence on $k_{\chi}$, $\xi$, $m_{\chi}$, and $\Delta t$. In the following, we present the numerical results for Eq. (6.15) in Table 1. From our numerical analysis, by varying $k_{\chi}$ and $m_{\chi}$ with fixed $\xi$, it appears that $\Delta t$ is insensitive to the chosen parameters. The duration required to obtain a sufficient amount of PBHs is found to be $\sim 10^{7}M_{p}^{-1}$. This value is consistent with Fig. 2, which shows that the effective growth of $\chi$ production occurs after $t>10^{6}M_{p}^{-1}$. Note that in our numerical calculation, we keep $m_{\chi},k_{\chi}\ll\xi M^{2}$. In this case, the PBH constraint may not jeopardize our model. Moreover, even if $\chi$ grows larger, at some point it cannot collapse into PBHs due to the rapidly expanding universe. Thus, PBHs are suppressed beyond a certain stage [green2001primordial].
7 Reheating scenario
In this section, we will discuss the possible reheating scenario in this model. It is noted that the reheating stage is the condition where the cosmological model can be closely linked to SM particle physics. However, it turns out that calculating the reheating temperature in 3 is not straightforward, but using the e-fold connection disregards the actual mechanism. We will see in the following subsection that several mechanisms can appear, and we will discuss which mechanism is the most favored.
7.1 The Perturbative decay of the scalaron to scalar $\chi$
The most traditional depiction of the reheating temperature is based on the simple perturbative decay of the inflaton into light and relativistic particles. It is assumed that the energy of these light particles is converted into radiation. If we consider the scalaron decaying to $\chi$ via the Lagrangian
$\sqrt{\frac{3}{2}}\frac{\phi}{M_{p}}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi,$ (7.1)
the decay rate, within approximation $m_{\chi}\ll M$, can be written as
$\Gamma(\phi\rightarrow\chi\chi)\simeq\frac{3}{128\pi}\frac{M^{3}}{M_{p}^{2}},$ (7.2)
and the reheating temperature can be easily evaluated as
$T_{\text{reh}}=\left(\frac{90}{g^{*}\pi^{2}}\right)^{1/4}\sqrt{M_{p}\Gamma},$ (7.3)
where $g^{*}\sim 100$ is the SM degrees of freedom. As $M$ is well-constrained by ACT, it is clear that the reheating temperature can be obtained as
$T_{\text{reh}}\approx 3.14\times 10^{9}\text{GeV}.$ (7.4)
However, this value is 5 orders of magnitude larger than our prediction after applying the ACT constraint. In this case, the reheating temperature from this perturbative mode is strongly discouraged. One should note that this value could be enhanced by BEC effects, which may increase the reheating temperature.
7.2 The Perturbative decay of the scalaron to fermion $\psi$
As previously remarked, the perturbative decay of the scalaron to $\chi$ is heavily suppressed by ACT results. In that case, the interaction between the scalaron and fermions is described by the Dirac Lagrangian in the Einstein frame, which reads
$\mathcal{L}_{\text{Dirac}}=e^{-3\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m_{\psi})\psi,$ (7.5)
where $\bar{\psi}$ and $\psi$ are antifermion–fermion pairs, $\gamma^{\mu}$ is the Dirac gamma matrix, and $m_{\psi}$ is the mass of the fermions. Due to the kinetic term of the Dirac Lagrangian being on-shell, the interaction term of the scalaron with $\bar{\psi}\psi$ is controlled by
$3\sqrt{\frac{2}{3}}\frac{m_{\psi}}{M_{p}}\phi\bar{\psi}\psi.$ (7.6)
Straightforwardly, we obtain the decay rate as
$\Gamma(\phi\rightarrow\bar{\psi}\psi)\simeq\frac{3}{4\pi}\frac{m_{\psi}^{2}M}{M_{p}^{2}},$ (7.7)
and the reheating temperature as high as
$T_{\text{reh}}\simeq 8.07\times 10^{-4}m_{\psi}.$ (7.8)
By using the ACT reheating temperature constraint of $10^{4}$ GeV, $m_{\psi}\sim 10^{8}$ GeV could satisfy the prediction. Generally speaking, this mechanism is favored but we will see later that at some point, a certain mechanism is more favored.
7.3 Remarks on the perturbative mode
In our discussion about the perturbative reheating by scalar $\chi$ and fermion $\psi$, it is shown that the scalar mode is disfavored by ACT results as it produces a high reheating temperature. Furthermore, as it is enhanced by BEC, the reheating temperature from this mode could be larger than $10^{9}$ GeV. On the other hand, reheating via fermions is more favorable, as it naturally allows for lower temperatures and links the connection more into Standard Model particles due to their chiral nature [becker2024confronting].
In Ref. [allahverdi2010reheating], it is shown that both reheating mechanisms are rather oversimplified, as they neglect the violation of the fluctuation-dissipation theorem (see Refs. [hu1995fluctuation, berera2005absence, gleiser1994microphysical]). Additionally, these mechanisms overlook the effects of preheating and can only be significant if preheating is inefficient. This regime slowly transitions to a radiation-dominated regime, and the perturbative decay channel opens only gradually. It is implied that perturbative reheating via bosons is strongly disfavored due to the high reheating temperature, which contradicts our low reheating temperature prediction. Also, during the matter-dominated regime, all dominant particle production is mostly heavy non-relativistic [kofman1997towards]. On the other hand, fermionic reheating suffers not only from the same problem but also from the Pauli exclusion principle, which further suppresses their production. Even though this mode is more favorable than the bosonic one, its reheating scenario is still not convincing.
In addition, Ref. [risdianto2025second], which discusses inflation with a non-minimal coupling model, shows that perturbative reheating is strongly disfavored. This is because achieving a favorable reheating temperature requires violating unitarity.
7.4 Reheating temperature by preheating
As we learn from the perturbative reheating, it has some issues at certain points. Thus, the reheating temperature should be determined from the burst of particle production resulting from resonance. Our chosen mechanism is as follows: Due to the nature of the matter-dominated regime, it is expected that the reheating temperature will be evaluated at the end of the preheating stage. However, the reheating temperature should not be evaluated from the ordinary perturbative decay of either bosons or fermions, as discussed in the previous subsections. Instead, the reheating temperature should be evaluated from the last burst of scalaron oscillation that produces scalar $\chi$ at the end of the preheating stage [^7] This was discussed in Ref. [risdianto2025second]). Although in this reference, it used the beginning of the second stage of preheating.. The produced scalar $\chi$ will immediately decay into relativistic fermions $\psi$ to reheat the universe [^8] We choose the bosons $\chi$ as the product at the zero crossing of the scalaron due to the abundance of bosons over fermions during preheating. Fermions $\psi$ are chosen as the decay product of $\chi$ because they are more directly related to SM particles due to their chiral nature.. Intuitively, we can assume that during this time, the remaining energy density of the inflaton is very small. Thus, a low reheating temperature is expected.
To calculate this reheating temperature, it is written as
$T_{\text{reh}}=\left(\frac{30}{g^{*}\pi^{2}}\delta\rho\right)^{1/4},$ (7.9)
where $\delta\rho$ is the energy density produced by the single crossing during the ’last’ preheating. The $\delta\rho$ can be evaluated as [risdianto2025second, hashimoto2021inflation]
$\delta\rho=\int^{\pi M^{-1}}_{0}dt\hskip 2.84526pt\Gamma_{\chi\rightarrow\bar{\psi}\psi}n_{\chi}\bar{m}_{\chi}e^{-\int^{t}_{0}\Gamma_{\chi\rightarrow\bar{\psi}\psi}dt^{\prime}},$ (7.10)
where
$n_{\chi}=\int\frac{d^{3}k_{\chi}}{(2\pi)^{3}}e^{2\mu_{p}Mt}\simeq\frac{k_{\chi}^{3}}{6\pi^{2}}$ (7.11)
is the number density of $\chi$. We set $\mu_{p}\simeq 0$ since during this time the preheating is almost ended. The mass $\bar{m}_{\chi}$ is estimated from the Eq. (4.5) as
$|\bar{m}_{\chi}|=\sqrt{\sqrt{6}\frac{\xi M^{2}\tilde{\phi}}{M_{p}}\sin(Mt)}.$ (7.12)
To complete Eq. (7.10) we should consider that there is an interaction between $\chi$ and $\psi$ via
$y\chi\bar{\psi}\psi,$ (7.13)
where $y$ is the Yukawa coupling between $\chi$ and $\bar{\psi}\psi$. The decay rate for this channel is approximately given by,
$\Gamma_{\chi\rightarrow\bar{\psi}\psi}\simeq\frac{y^{2}m_{\chi}}{8\pi}.$ (7.14)
With these, we can solve the Eq. (7.10) as
$\delta\rho\simeq 3.71\times 10^{-6}y^{2}k_{\chi}^{3}m_{\chi}\sqrt{\xi}$ (7.15)
Straightforwardly, the reheating temperature can be estimated as
$T_{\text{reh}}\simeq 1.83\times 10^{-2}\left(y^{2}k_{\chi}^{3}m_{\chi}\sqrt{\xi}\right)^{1/4}\quad\text{GeV}.$ (7.16)
The reheating temperature is now shown to be dependent on several parameters. In short, one can see the Table. 2. It is shown that low $y$, $k_{\chi}$, and $m_{\chi}$ are required to achieve the favored reheating temperature of $\sim 10^{4}$ GeV.
$y$ $k_{\chi}$ $m_{\chi}$ $T_{\text{reh}}$
$10^{-1}$ $10^{-9}M_{p}$ $10^{-16}M_{p}$ $4.46\times 10^{5}$ GeV
$10^{-2}$ $10^{-9}M_{p}$ $10^{-16}M_{p}$ $1.41\times 10^{5}$ GeV
$^{-3} 1010$$^{-10}M_{p}$ $10^{-12}M_{p}$ $4.46\times 10^{4}$ GeV
$^{-4} 1010$$^{-10}M_{p}$ $10^{-12}M_{p}$ $1.41\times 10^{4}$ GeV
$10^{-4}$ $10^{-9.5}M_{p}$ $10^{-12}M_{p}$ $3.34\times 10^{4}$ GeV
$10^{-4}$ $10^{-9}M_{p}$ $10^{-12}M_{p}$ $7.92\times 10^{4}$ GeV
$10^{-4}$ $10^{-8.5}M_{p}$ $10^{-12}M_{p}$ $1.88\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-12}M_{p}$ $4.46\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-13}M_{p}$ $2.51\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-14}M_{p}$ $1.4\times 10^{5}$ GeV
$10^{-4}$ $10^{-8}M_{p}$ $10^{-15}M_{p}$ $7.92\times 10^{4}$ GeV
Table 2: The reheating temperature with varying $y$, $k_{\chi}$ and $m_{\chi}$. We set $\xi=1$ and adjust the parameters to be in the range of $T_{\text{reh}}$ of $10^{4}\sim 10^{5}$ GeV.
Considering the chosen parameters, we need to clarify some favored values. The coupling $y$ should be small to suppress instant reheating (see [bezrukov2009initial]). For $m_{\chi}$, this lies close to the electroweak scale and should be higher. A model with this mass range favors linkage to the SM. However, as $\chi$ and fermion $\psi$ have bare masses, and their mass does not come from the symmetry breaking, they should not belong to the SM particles. [^9] As it is noted that the SM particles are massless at higher than the electroweak scale Thus, for safety, $m_{\chi}\gg m_{\psi}\gg 10^{-16}M_{p}$. In conclusion, both $\chi$ and $\psi$ masses should lie above the electroweak scale, as neither fields behave like an SM particle. Note that $y$ could be much smaller. Thus, all other parameters could be set to higher values, $k_{\chi}\gg 10^{-8}M_{p}$ and $m_{\chi}\gg 10^{-16}M_{p}$. In this model, it seems that the reheating temperature is weakly constrained by parameters $y$, $k_{\chi}$, and $m_{\chi}$.
Lastly, Table 2 presents the preferred parameter values that accommodate the constrained reheating temperature. From the table, it can be seen that a reheating temperature of approximately $10^{4}$ GeV can be achieved. These parameters can also be adjusted to yield a higher reheating temperature (e.g., $10^{8}$ GeV), which is consistent with the previous Planck results. However, when applying the bound $n_{s}=0.9743\pm 0.0034$, the expected reheating temperature decreases to only a few hundred MeV. Therefore, our proposed mechanism fails under this condition and requires further investigation in future work [Risdianto2026PreheatingStarobinsky].
8 Conclusion
In this paper, we reinvestigate Starobinsky inflation, particularly the preheating stage, following the release of the ACT results. These results suggest a significant increase in the e-fold number ($N\simeq 75$) and constrain the reheating temperature to $T_{\text{reh}}\sim 10^{4}$ GeV if we adopt the center value with $68\%$ of the confidence level. Consequently, the preheating and reheating mechanisms must be modified to align with these new results. Such a low reheating temperature affects the entire preheating stage, not just the reheating mechanism. In addition, the direct perturbative decay from the scalaron to scalar $\chi$ is strongly disfavored if the preheating mechanism is dominant and the high reheating temperature is implied. On the other hand, reheating via the decay of the scalaron to fermions $\psi$ is also disfavored if the preheating stage is dominant, although less severely, due to the possibility of achieving a lower reheating temperature and its chirality nature.
For the efficient preheating to occur, a certain initial condition of $\chi$ needs to exist at the start of preheating. This initial condition is believed to have existed because $\chi$ is a spectator field during inflation. Surprisingly, this burst of resonance particles does not occur during the inflaton zero-crossing, as many inflationary models do, but rather at the moment when tachyonic preheating takes place. However, tachyonic preheating itself does not seem to be efficient enough to drain the inflaton’s energy density at once. This is a rather peculiar condition compared with many other models of inflation.
Ultimately, the reheating mechanism is realized through a specific preheating process. During resonance, the oscillating scalaron produces the scalar $\chi$. While fermion production is also possible, it is heavily suppressed due to the Pauli exclusion principle. During the preheating stage, perturbative decay of the scalaron and $\chi$ occurs but proceeds very slowly and does not contribute adequately to the reheating temperature. At the end of the preheating stage, the final burst of particle production from the scalaron to $\chi$, followed by the immediate decay of $\chi$ into fermions $\psi$, becomes the sole contributor to the reheating temperature. The reheating mechanism via this mode is found to be consistent with our chosen parameters.
In the following, we have also considered the favored parameters in our model based on our discussion in the previous sections. In our model, $y$ should be small to avoid instant reheating, which would lead to a high reheating temperature. The non-minimal coupling $\xi$ should be treated carefully, since the preheating could not be efficient enough if $\xi\simeq 1$ and should not be greater than $10$ to avoid the unitarity issues. Also, $m_{\chi}$ should be lower than $\frac{1}{2}M$ to ensure that the scalaron decays completely and higher than the electroweak scale ($\gtrsim 10^{3}$ GeV $\approx 10^{-15}M_{p}$), as $\chi$ is shown to be different to the SM particles to preserve the viability of our model [^10] If $m_{\chi}$ is at the electroweak scale and not the Higgs field, it may have already been observed. Also, the mass is not coming from the symmetry breaking in which should not belong to SM particles in the first place.. The mass $m_{\psi}$ should be much lower than $m_{\chi}$ to ensure a successful reheating scenario [^11] As $\psi$ is supposed to reheat the universe, it should be relativistic. Thus, $m_{\psi}\ll m_{\chi}$ is preferred.. One should note that, as $\chi$ in the early stage of preheating is a heavy, non-relativistic particle, it is implied that $k_{\chi}$ should be much lower than $\xi M^{2}$. Finally, Table 3 provides the appropriate values of the parameters. It is also shown that the chosen parameters could allow both $\chi$ and $\psi$ to be detected in future colliders.
In addition, our preheating mechanism successfully describes scenarios with a lower reheating temperature but fails when the temperature drops below $10^{3}$ GeV. The uncertainty in the ACT result, $n_{s}=0.9743+0.0034$, could lead to a reheating temperature on the order of below the GeV scale (approximately $10^{2}$ MeV), under which our preheating mechanism no longer works. This issue will be addressed in future work.
Parameter value Note
$y$ $\ll 10^{-3}$ avoid instant preheating
$\chi(0)$ $\sim 10^{-3}M_{p}$ efficient preheating
$\xi$ $\approx 10$ constraint: unitarity issue (upper bound), and efficient preheating (lower bound)
$m_{\chi}$ $10^{-6}\sim 10^{-14}M_{p}$ $\frac{1}{2}M$ (no inflaton’s remnant condition) and $\gg$ electroweak scale
$m_{\psi}$ $10^{-7}\sim 10^{-15}M_{p}$ $\ll m_{\chi}$ and $\gg$ electroweak scale
$k_{\chi}^{2}$ $\ll 10^{-11}M_{p}^{2}$ $\ll\xi M^{2}$ (efficient preheating at the end of inflation)
Table 3: The viable values of the parameters in our model.